OpenStudy (anonymous):

I don't know how to do this problem? :( 3.sqrt(18x^7y^6) I am supposed to simplify it

OpenStudy (amistre64):

divide the exponents in the sqrt by 2

OpenStudy (amistre64):

\[\sqrt{a^n}=a^{n/2}\]

OpenStudy (anonymous):

?? 7/2 and 3??

OpenStudy (amistre64):

thats a good start yes what is 7/2 as a mixed number?

OpenStudy (amistre64):

one approach could be like this 3.sqrt(18x^7y^6)\[3~~18^{1/2}x^{7/2}~y^{6/2}\] \[3~~18^{1/2}x^{3+1/2}~y^{3}\] \[3~~(9*2)^{1/2}x^{3+1/2}~y^{3}\] \[3~~9^{1/2}~2^{1/2}x^{3}~x^{1/2}~y^{3}\] \[3~3~x^{3}~y^{3}~(2x)^{1/2}\] \[9x^{3}y^{3}\sqrt{2x}\]

OpenStudy (anonymous):

4 3/4??

OpenStudy (amistre64):

7/2 = 3 + 1/2

OpenStudy (anonymous):

wouldn't the 9x^3y^3sqrt(2x) simplify to 3x^3y^3sqrt(2x) ??

OpenStudy (amistre64):

simplify? no. as a rule of thumb, the fewer elements you have, the simpler it is

OpenStudy (amistre64):

9 is one element 3 3 is two elements

OpenStudy (amistre64):

i might have read that a little off tho

OpenStudy (amistre64):

would 9 x^3 y^3 sqrt(2x) simplify to 3 x^3 y^3 sqrt(2x) no, since 9 is not 3

OpenStudy (anonymous):

Okay, but 9 x^3 y^3 sqrt(2x) is not one of my options, and 3 x^3 y^3 sqrt(2x) is.. I tried figuring this out on my own, but I cant seem to get it

OpenStudy (amistre64):

make sure youve posted the correct information:\[3\sqrt{x^7y^6}\]right?

OpenStudy (amistre64):

forgot an 18 under there by accident lol

OpenStudy (amistre64):

or is that 3. in front a label of the thrid problem?

OpenStudy (anonymous):

No, I made a mistake, I apologies. There is no 3. It is just sqrt18x^7y^6

OpenStudy (amistre64):

then yes, 3 x^3 y^3 sqrt(2x) should suffice :)

OpenStudy (anonymous):

Thanks so much :) Do you think you could help me through another one?

OpenStudy (amistre64):

i can try

OpenStudy (anonymous):

sqrt(3) - sqrt(6) / sqrt(3) + (sqrt(6)

OpenStudy (amistre64):

looks to be a conjugate operation

OpenStudy (amistre64):

does it say to rationalize the top or bottom?

OpenStudy (anonymous):

It says to give the simplest form of the equation @amistre64